OpenStudy (anonymous):

find the distance between the points (9,9) and (-7,-7)

pooja195 (pooja195):

Use the distance formula \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

pooja195 (pooja195):

(x1,y1)(x2,y2)

OpenStudy (anonymous):

so how would I enter the values

pooja195 (pooja195):

\[\huge d=\sqrt{(-7-9)^2+(-7-9)^2}\]

OpenStudy (anonymous):

ok

pooja195 (pooja195):

Yea....now solve that :)

OpenStudy (anonymous):

why dont u solve it for him... Qualified helper O_O

pooja195 (pooja195):

I dont give the answers.

pooja195 (pooja195):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Really? That's all I look for 90% of the time

pooja195 (pooja195):

@drkid15 have you solved it?

OpenStudy (anonymous):

wouldnt it just be (-7 - 9) + (-7 - 9)? cause its sqrted?

OpenStudy (anonymous):

d = -32?

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

im guessing so

OpenStudy (anonymous):

is ur question multiple choice?

OpenStudy (anonymous):

no

pooja195 (pooja195):

Why not try solving it yourself?

pooja195 (pooja195):

If you need help i can GUIDE you.

OpenStudy (anonymous):

idk how

OpenStudy (anonymous):

Im, 99% sure its -32, but I am blond... so I might be stupid

OpenStudy (anonymous):

can u just guide me plz

OpenStudy (anonymous):

Alright, so the equation is \[D = \sqrt{(x2-x1)^{2} + (y2 - y1)^{2}}\] right?

OpenStudy (anonymous):

D is the distance between 9,9 and -7,-7

OpenStudy (anonymous):

both of those sets have an x and a y its like this (x,y)

OpenStudy (anonymous):

but since theres 2 sets, were gonna label it like (x1,y1) and (x2,y2)

OpenStudy (anonymous):

Are you paying attention Drkid15?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

For the equations we have 9,9 as the first pair, so it takes over the x1 and y1

OpenStudy (anonymous):

can u show me what it looks like when u enter the values

OpenStudy (anonymous):

\[D = \sqrt{(9 - (-7))^{2} + (9 - (-7)^{2}}\]

OpenStudy (anonymous):

By the way yes, the -32

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

o wait no